OpenStudy (anonymous):

A circle has a circumference of 14 ft. Its area is halved. Describe the effect on the circumference of the circle.

OpenStudy (anonymous):

first find radius 2 pi r=14 r=7/Pi plug that into area formula pi (7/pi)^2= 49/Pi then half that and see what happens to radious 49/pi*.5=pi r^2 sfr 4.949747468/Pi now plug it back into circumference 2*pi*(4.949747468/Pi)= 9.899494936 so original is 14 when you half the area you get 9.899 14/9.899=1.414 shrinkage

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

de nada :)